NEAR INFRARED LASER SPECTROSCOPY OF NiI

W. S. TAM, JIANJUN YE and \underline{A. S-C. CHEUNG}, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong..

Laser induced fluorescence spectrum of NiI in the near infrared region 714 - 770 nm has been recorded and analyzed. Eleven bands belonging to four electronic transition systems have been observed: the (0,0), (1,0), and (2,0) bands of [13.3] 2Σ^{+} - A ^{2}Π_{3/2} system; the (0,0), (1,1) and (0,1) bands of [13.9] ^{2}Π_{3/2} - X ^{2}Δ_{5/2} system; the (0,0) and (1,0) bands of [13.9] 2Π_{3/2} - A ^{2}Π_{3/2} system; and the (0,0), (1,1) and (0,1) bands of the [14.0] ^{2}Φ_{7/2} - X ^{2}Δ_{5/2} system. Spectra of isotopic molecules confirmed the vibrational quantum number assignment of the observed bands. Least squares fit of observed rotational lines yielded accurate molecular constants for the [13.3] ^{2}Σ^{+}, [14.0] ^{2}Φ_{7/2} and the A ^{2}Π_{3/2} electronic states of NiI. Hot band transition from the v = 1 level of the X ^{2}Δ_{5/2} state gives ΔG_{1/2} = 276.67 cm^{−1}. The A ^{2}Π_{3/2} state was determined to be only 163.85 cm^{−1} above the X ^{2}Δ_{5/2} state.